A random sample of 60 students is chosen from a large population with an average diastolic blood pressure of 100 and a standard deviation of 36. What is the probability that the mean diastolic blood pressure of these 60 students is less than 110?


1
Expert's answer
2022-03-14T18:28:49-0400

"P(\bar X<110)=P(Z<\frac{110-100}{\frac{36}{\sqrt{60}}})=P(Z<2.15)=0.9842."